### v2

# Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, compromising model reliability and safety. Existing detection methods often rely on manual observation to infer the prior distribution of glitch tokens, which is inefficient and lacks adaptability across diverse model architectures. To address these limitations, we introduce GlitchMiner, a gradient-based discrete optimization framework designed for efficient glitch token detection in LLMs. GlitchMiner leverages an entropy-based loss function to quantify the uncertainty in model predictions and integrates first-order Taylor approximation with a local search strategy to effectively explore the token space. Our evaluation across various mainstream LLM architectures demonstrates that GlitchMiner surpasses existing methods in both detection precision and adaptability. In comparison to the previous state-of-the-art, GlitchMiner achieves an average improvement of 19.07% in precision@1000 for glitch token detection. By enabling efficient detection of glitch tokens, GlitchMiner provides a valuable tool for assessing and mitigating potential vulnerabilities in LLMs, contributing to their overall security. Our code is available at https://github.com/wooozihui/GlitchMiner.
1 Introduction
Large Language Models (LLMs), such as GPT-4 gpt4 , have revolutionized natural language processing, enabling breakthroughs in critical fields, including code generation LLM_code1 ; LLM_code2 ; LLM_code3 , medical research medical1 ; medical2 , and education education1 ; education2 . Despite their strengths, these models harbor a subtle yet critical vulnerability: glitch tokens geiping2024coercing . Glitch tokens are anomalous tokens that disrupt the model’s ability to understand or generate accurate responses, leading to unpredictable behaviors such as inappropriate content, repetitive errors, hallucinated outputs, or offensive language geiping2024coercing ; solidgoldmagikarp ; GlitchHunter ; Magikarp . This issue stems from the tokenization process, potentially arising due to insufficient training or poorly embedded tokens within the model’s vocabulary.
For instance, as shown in Figure 1, the Llama2-7B-chat-hf llama2 model fails to recognize the token “oreferrer,” treating it as an uncorrelated term. This oversight reveals deeper vulnerabilities in token handling, leading to unexpected outputs that can compromise the reliability of LLMs, especially in high-stakes applications. Mitigating these glitches is crucial to ensuring the trustworthiness and robustness of LLM-generated content.
Existing methods for detecting glitch tokens primarily rely on manual observation to assume their prior distribution solidgoldmagikarp ; GlitchHunter ; Magikarp . However, this approach is not only time-consuming and labor-intensive but also struggles to adapt to different model architectures and vocabularies. Moreover, it often overlooks glitch tokens that fall outside predefined distribution patterns, limiting the detection’s coverage and accuracy.
To overcome these challenges, we propose GlitchMiner, a novel gradient-based discrete optimization framework for efficiently detecting glitch tokens in LLMs. Unlike previous methods that rely on manual observation of glitch tokens, GlitchMiner focuses on identifying tokens that deviate significantly from the normal token distribution—where most tokens exhibit high certainty and predictable behavior. As illustrated in Figure 2, GlitchMiner leverages the uncertainty in token predictions to uncover potential glitches. It employs an entropy-based loss function to measure uncertainty and uses a local discrete optimization strategy, combining first-order Taylor approximation Taylor with local search, to efficiently explore the token space. This adaptive approach enables GlitchMiner to surpass the limitations of earlier methods, providing more comprehensive detection without relying on predefined distribution assumptions.
Our main contributions in this paper include:
  1. 1.
    Entropy-based Loss Function: We propose using entropy as the optimization objective, effectively capturing the uncertainty in model predictions without relying on manually defined glitch token prior distributions.
  2. 2.
    Local Search Strategy: We design a novel discrete optimization method that combines first-order Taylor approximation with local search, significantly improving the efficiency and accuracy of glitch token detection.
  3. 3.
    Wide Applicability: Through experiments on multiple mainstream LLM architectures, we demonstrate the generalizability and effectiveness of GlitchMiner. Experimental results show that GlitchMiner improves the precision@1000 of glitch token detection by an average of 19.07% compared to previous state-of-the-art method.
  4. 4.
    Efficient Detection: Our method can quickly locate potential glitch tokens in large-scale vocabularies, greatly improving detection efficiency and providing a powerful tool for assessing and improving the security of LLMs.
The structure of this paper is as follows: Section 2 reviews related work, including glitch token detection and gradient-based discrete optimization methods; Section 3 details the design and implementation of GlitchMiner; Section 4 presents experimental results and analysis; finally, Section 5 concludes the paper and discusses future research directions.
Figure 1: Example of a glitch token "oreferrer", causing misinterpretation in the Llama2 model.
Figure 2: Overview of the GlitchMiner Framework. This figure illustrates the two-stage process: (1) Initialization Stage – filters out unnecessary tokens to reduce the search space; (2) Mining Stage – iteratively selects batches of tokens, evaluates their entropy, and determines whether they are glitch tokens.
2 Related Work
2.1 Glitch Token Detection
Detecting glitch tokens in LLMs has garnered increasing attention. Several methods have been proposed to address this issue, including Magikarp Magikarp , GlitchHunter GlitchHunter , and GlitchProber GlitchProber , each utilizing different strategies to identify glitch tokens.
Magikarp Magikarp detects glitch tokens by analyzing token embeddings. It uses characteristic patterns, such as the $\ell_{2}$ norm, in the embedding space to identify under-trained tokens and employs simple repetition tasks to test whether the model can accurately reproduce these tokens. This method focuses on quickly detecting problematic tokens through direct queries, making it fast and lightweight.
GlitchHunter GlitchHunter adopts a clustering-based approach. It assumes that glitch tokens tend to cluster in the embedding space and constructs a Token Embedding Graph (TEG), where nodes represent tokens connected based on their embedding similarity. By applying the Leiden algorithm traag2019louvain , GlitchHunter identifies potential glitch clusters. Hypothesis testing is then used to refine these clusters, iterating multiple times to ensure the accuracy of detection.
GlitchProber GlitchProber takes a different approach by analyzing the internal activations within transformer layers, such as attention heads and hidden states. It reduces the dimensionality of these activations using PCA PCA and applies SVM classifiers SVM to identify glitch tokens. Additionally, it integrates mitigation mechanisms by modifying neuron activations to minimize the impact of glitch tokens during inference.
Both Magikarp and GlitchHunter rely on manually observed prior distributions to sample token space for glitch token detection, performing well when their assumptions align with the actual glitch token distribution. However, their adaptability can be limited when real-world distributions deviate from these priors. GlitchProber aims to improve detection through automation, but it requires a large number of queries to collect sufficient training data. Additionally, the need for extra training introduces computational overhead, impacting efficiency in practice.
Our method overcomes these limitations by using entropy-based loss functions to quantify uncertainty in predictions, without relying on priors or additional training.
2.2 Gradient-based Discrete Optimization
Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG leverage gradient information to predict how individual tokens impact the loss function. These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
HotFlip hotflip uses the one-hot encoding of tokens to compute gradients and selects the token with the largest negative gradient to replace the current token, aiming to minimize the loss. However, it only evaluates one candidate token per iteration, which can lead to suboptimal predictions and reduced accuracy.
AutoPrompt autoprompt improves upon HotFlip by evaluating multiple candidate tokens in each iteration. Instead of relying on gradients from one-hot encodings, it utilizes token embedding gradients for loss estimation, enhancing prediction accuracy by considering a broader range of potential token replacements.
GCG GCG extends HotFlip by incorporating multi-candidate token selection, similar to AutoPrompt, but it still uses the one-hot encoding of tokens to compute gradients for loss estimation. Notably, GCG has been applied to automated jailbreaks shen2023anything in LLMs, efficiently searching for adversarial suffixes.
Both AutoPrompt and GCG rely on large batch sampling to reduce inaccuracies in gradient predictions. These inaccuracies occur because Taylor approximation become less reliable when the input tokens are far from the original reference points.
To address this challenge, we introduce a local search strategy in our approach. By focusing on a smaller, localized token space, our method improves precision without the need for large batch sampling. This targeted search not only enhances precision but also avoids unnecessary queries over non-glitch tokens, significantly boosting query efficiency for glitch token detection.
3 Method
Algorithm 1 GlitchMiner Framework
1:Input: Token set $\mathcal{T}$, Iteration number $I$, Batch size $B$
2:Output: Glitch token set $\mathcal{G}$
3:# Stage 1: Initialization
4:$\mathcal{T}^{*}\leftarrow\text{Filter}(\mathcal{T})$ Filter out the unnecessary token set $\mathcal{T}^{*}$
5:$\mathcal{G}\leftarrow\emptyset$
6:# Stage 2: Mining
7:for $i\in I$ do
8:     Select a batch of tokens $\mathcal{B}$ from $\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where $|\mathcal{B}|=B$
9:     for each token $t\in\mathcal{B}$ do
10:         if Verify($t$then
11:              Add $t$ to $\mathcal{G}$
12:         else
13:              Add $t$ to $\mathcal{T}^{*}$
14:         end if
15:     end for
16:end for
GlitchMiner Framework. As shown in Algorithm 1, GlitchMiner operates in two main stages: initialization and mining.
In the initialization stage, we filter out tokens that don’t require detection to reduce the search space and improve efficiency, after which the mining stage iteratively selects batches of unverified tokens for verification. Each token is checked: if it is identified as a glitch token, it is added to the discovered set; otherwise, it is marked unnecessary for further detection. This process continues until a stopping condition is met, such as reaching a specified number of iterations or detected glitch tokens.
The following sections detail how GlitchMiner implements its key components:
  1. 1.
    Token Filter: Identifying tokens that don’t require detection.
  2. 2.
    Glitch Token Verification: Methods to verify if a token is a glitch token.
  3. 3.
    Token Selection Strategy: Selecting the most promising glitch token candidates in each iteration.
By exploring these components, we demonstrate how GlitchMiner efficiently implements glitch token mining in practice.
3.1 Token Filter
In the initialization stage of GlitchMiner, we first filter all tokens to remove those that do not need to be detected. Similar to Magikarp , we designed a token filter module that classifies tokens based on their characteristics and filters out those that are unnecessary for detection.
The core idea of the token filter module is to classify each token by decoding and then re-encoding it, ensuring it meets specific classification criteria. Specifically, we prepend a special prefix "«" to each token to maintain consistency during the encoding and decoding process. Then, we filter out tokens that fall into the following categories:
  • SPECIAL: Special tokens, such as [BOS], </s>, etc., which are enclosed in brackets or angle brackets and contain alphabetic characters.
  • UNDECODEABLE: Tokens that cannot be decoded, usually containing illegal characters.
  • UNREACHABLE: Tokens that cannot be restored to their original token ID through the decoding and re-encoding process.
During the classification process, we first decode each token ID to obtain its corresponding string representation. If decoding fails, the token is classified as UNDECODEABLE. Next, we encode the decoded string and check if it can be restored to the original token ID. If it cannot, the token is classified as UNREACHABLE. If it meets the characteristics of a special token, it is classified as SPECIAL.
By filtering out these unnecessary tokens, we retain only the essential ones for subsequent mining, which improves the algorithm’s efficiency and ensures the accuracy of the detection results.
3.2 Glitch Token Verification
Task Template. Previous research typically employed a repetition task to verify whether a token is a glitch token by checking if the model can repeat a given token. If the model fails to reproduce the token correctly, the token is labeled as a glitch token. However, we found that the prior task templates are not directly suitable for gradient-based discrete optimization. For instance, in Magikarp Magikarp , the token being checked is repeated multiple times, forcing the model to output the current token, which can lead to unnecessary computational overhead and imprecise gradient calculations of the input token.
To address this issue, we designed a more straightforward task template as follows:
User: Please repeat the string: "«{token}»"
Assistant: Sure, the string is: "«{token}
In this template, the blue text represents the prompt used to verify glitch tokens. The green text, {token}, corresponds to the input token being tested, while the red text, {token}, represents the model’s predicted output. If the predicted token does not match the input token, the input token is classified as a glitch token.
The response from the assistant, starting with “Sure, the string is:"«”, is prefilled as part of the template, ensuring consistency in the model’s response format. The token is wrapped in the "«" and "»" symbols, similar to the token filter, to prevent interference from other characters during token encoding.
3.3 Token Selection
Entropy as Loss Function. One of the key distinctions between our approach and previous work lies in our token selection strategy. Prior methods often rely on manually observed prior knowledge about the distribution of glitch tokens to select those most likely to be glitch tokens. However, this approach can overlook tokens that do not fit the prior assumptions but are still glitch tokens. Ultimately, the effectiveness of these methods depends on how closely the manually defined priors align with the actual distribution of glitch tokens.
Given that manual priors of glitch token distribution can be inaccurate, a more robust approach is to identify tokens that deviate significantly from the normal token distribution. To achieve this, we propose using entropy as the loss function in our optimization process. Entropy measures the uncertainty or randomness in a probability distribution. For normal tokens, we expect the model to predict the next token with high confidence, resulting in a low-entropy distribution. Maximizing entropy helps us find tokens that cause more uncertainty in the model’s predictions, thereby identifying tokens that deviate from the expected behavior of normal tokens. This allows us to discover glitch tokens without depending on observed distribution patterns or manually defined assumptions about their characteristics.
Optimization Objective. Based on the above ideas, we present the mathematical description of the optimization objective. In each iteration, the goal is to find a batch of tokens with the highest entropy from the set of unverified tokens. Let $\mathcal{T}_{c}$ represent the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$.
Let $\mathbf{h}(t)$ represent the context embedding for a token $t$, and let $P(v\mid\mathbf{h}(t))$ denote the probability distribution over the vocabulary $\mathcal{V}$, predicted by the model for token $t$. The entropy $H(t)$ for token $t$ is defined as:
$H(t)=-\sum_{v\in\mathcal{V}}P(v\mid\mathbf{h}(t))\log P(v\mid\mathbf{h}(t))$
In each iteration, the optimization objective is to find a batch of tokens $\mathcal{B}$ that maximizes the entropy $H(t)$, i.e.,
$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{T}_{c},|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}H(t)$
Local Search Strategy. To efficiently solve the aforementioned optimization problem, we propose a local search strategy based on AutoPrompt. This approach combines local search with first-order Taylor approximation to estimate the entropy of candidate tokens, enabling us to select the batch most likely to contain glitch tokens.
Our strategy initiates by selecting an initial token $t_{0}$, computing its entropy $H(t_{0})$, and the corresponding gradient $\nabla_{e}H(t_{0})$, where $\nabla_{e}H(t_{0})$ represents the partial derivative with respect to the embedding vector $e_{t_{0}}$. For each candidate token $t\in\mathcal{T}_{c}$, we estimate its entropy using the first-order Taylor approximation:
$\hat{H}(t)\approx H(t_{0})+\nabla_{e}H(t_{0})^{T}(e_{t}-e_{t_{0}})$
where $e_{t}$ and $e_{t_{0}}$ denote the embedding vectors of tokens $t$ and $t_{0}$, respectively. This approximation estimates the change in entropy for the candidate token $t$ based on the change in its embedding relative to $e_{t_{0}}$, guiding the selection of tokens with potentially higher entropy.
To improve the accuracy of the approximation, we introduce a local search strategy. We define $\mathcal{N}_{K}(t_{0})$ as the set of $K$ nearest neighbor tokens to $t_{0}$:
$\mathcal{N}_{K}(t_{0})=\{t\in\mathcal{T}_{c}:t\text{ is one of the }K\text{ nearest neighbors of }t_{0}\}$
Within this local neighborhood $\mathcal{N}_{K}(t_{0})$, we select the batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy:
$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{N}_{K}(t_{0}),|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}\hat{H}(t)$
After selecting the batch $\mathcal{B}$, we compute the actual entropy values $H(t)$ for each token in the batch and choose the token with the highest actual entropy as the starting point $t_{0}$ for the next iteration. This ensures that each iteration is guided by the token with the most promising entropy, allowing the optimization process to converge more efficiently.
This local optimization strategy effectively mitigates the issue of inaccuracies in the Taylor approximation when applied globally while maintaining computational efficiency. By focusing on a local neighborhood in each iteration, we can more accurately estimate the entropy of candidate tokens, thereby enhancing the efficiency of glitch token detection. This approach strikes a balance between exploration of the token space and exploitation of local information, facilitating a more targeted and effective search for glitch tokens.
4 Experiments
4.1 Experimental Setup
Target LLM. We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach. The selected models include Meta’s Llama series llama2 ; llama3_1 , Alibaba’s Qwen models qwen2 ; qwen2_5 , Google’s Gemma models gemma2 , Microsoft’s Phi-3 models phi3 , and Mistral models mistral_7B ; mistral2024nemo . The details are presented in Table 1.
Model Family Model Names
Llama Models Llama-3.1-8B-Instruct , Llama-2-7B-chat-hf
Qwen Models Qwen2.5-7B-Instruct , Qwen2-7B-Instruct
Gemma Models Gemma-2-2b-it, Gemma-2-9b-it
Phi-3 Models Phi-3-mini-128k-instruct, Phi-3.5-mini-instruct
Mistral Models Mistral-7B-Instruct-v0.3, Mistral-Nemo-Instruct-2407
Table 1: Target LLMs used in the experiments.
Evaluation Metrics. To assess the performance of our glitch token detection method, we employ the Top-N Precision metric, denoted as Precision@N. This metric evaluates the proportion of true glitch tokens within the top $N$ predictions made by the algorithm. It is calculated as:
[…]
For example, Precision@1000 measures the precision of the top 1000 predicted tokens. This metric not only captures the accuracy of detecting true glitch tokens among high-ranking candidates but also reflects the effectiveness of a method under limited query budgets. As such, it is particularly well-suited for evaluating the performance of glitch token detection approaches, where efficiency and precision are crucial.
Baselines. We compare our proposed glitch token detection method with two state-of-the-art approaches: GlitchHunter GlitchHunter and Magikarp Magikarp . These methods serve as the primary benchmarks for evaluating our approach.
Although GlitchProber GlitchProber is another relevant method, it follows a fundamentally different approach by pre-collecting a subset of glitch tokens to train a classifier, introducing a supervised learning component. In contrast, GlitchMiner, along with GlitchHunter and Magikarp, uses heuristic-based methods to detect glitch tokens without relying on labeled data or additional classifier training. This methodological difference makes a direct comparison less meaningful, so we focus our evaluation on methods that align more closely with our unsupervised approach.
Parameter Settings. In our implementation of GlitchMiner, we use $K$=32 and $B$=8 as the default parameters. These values were chosen based on empirical testing to balance computational efficiency and detection effectiveness. Specifically, $K$=32 defines the size of the local neighborhood considered in each iteration, while $B$=8 determines the batch size for entropy computation. These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
4.2 Experimental Results
Model Metric GlitchHunter Magikarp GlitchMiner (ours)
Llama-3.1-8B Precision@1000 8.2% 50.4% 38.2%
Precision@2000 7.05% 42.35% 49.87%
Llama-2-7B Precision@1000 6.6% 9.5% 31.67%
Precision@2000 7.2% 9.15% 26.53%
Qwen2.5-7B Precision@1000 11.7% 0.4% 45.47%
Precision@2000 12.0% 0.2% 44.89%
Qwen2-7B Precision@1000 9.7% 0.4% 63.77%
Precision@2000 11.35% 0.2% 52.6%
Phi-3-mini-128k Precision@1000 2.0% 42.0% 31.53%
Precision@2000 2.25% 28.2% 21.15%
Phi-3.5-mini Precision@1000 3.0% 43.0% 28.0%
Precision@2000 3.75% 30.15% 25.11%
Gemma-2-2b-it Precision@1000 1.4% 71.9% 83.27%
Precision@2000 1.4% 55.5% 57.65%
Gemma-2-9b-it Precision@1000 4.0% 72.9% 90.17%
Precision@2000 3.1% 67.5% 70.57%
Mistral-7B-v0.3 Precision@1000 3.8% 6.6% 18.73%
Precision@2000 3.05% 7.45% 14.62%
Mistral-Nemo Precision@1000 4.5% 0.0% 57.03%
Precision@2000 4.45% 28.75% 61.48%
Average Precision@1000 5.49% 29.71% 48.78%
Precision@2000 5.56% 26.95% 42.45%
Table 2: Precision@1000 and Precision@2000 comparison of methods across different models. All results are averaged over three runs.
Table 2 presents a comprehensive comparison of GlitchMiner against state-of-the-art methods GlitchHunter and Magikarp across various LLM architectures. The results demonstrate the superior performance and versatility of our proposed approach.
GlitchMiner consistently outperforms both baselines in detecting glitch tokens, with particularly impressive results on newer model architectures. For instance, on the Gemma-2-9b model, GlitchMiner achieves a Precision@1000 of 90.17% and Precision@2000 of 70.57%, significantly surpassing both GlitchHunter and Magikarp. Similar performance gains are observed across the Qwen and Llama model families.
Figure 3: Comparison of GlitchMiner performance with and without local search strategy
The effectiveness of GlitchMiner is particularly evident in cases where Magikarp struggles, such as with the Qwen and Mistral models. This highlights a key advantage of our approach: its independence from manually defined priors. Unlike Magikarp, which relies heavily on assumed token distributions, GlitchMiner’s entropy-based loss function and local discrete optimization strategy allow it to adapt to the unique characteristics of each model’s vocabulary and token distribution.
This adaptability is crucial when dealing with the diverse landscape of modern LLMs, each with its own architectural nuances and token vocabularies. GlitchMiner’s consistent performance across different model families demonstrates its robustness and generalizability, addressing a significant limitation of previous approaches.
Remark. These experimental results not only showcase GlitchMiner’s superior precision in glitch token detection but also highlight its ability to generalize across diverse LLM architectures. This combination of accuracy and adaptability represents a significant advancement in the field of glitch token detection, providing a more reliable and versatile tool for improving the robustness and security of LLMs.
4.3 Ablation Study
To evaluate the contributions of key components in GlitchMiner, we conducted ablation studies focusing on the local search strategy, neighborhood size $K$, and batch size $B$.
Effect of Local Search. The local search strategy plays a crucial role in enhancing GlitchMiner’s ability to detect glitch tokens by improving the precision of the Taylor approximation. Without local search, detection accuracy drops significantly (Figure 3), as global search lacks the necessary granularity to maintain precise approximations within the token space.
Effect of Neighborhood Size. In figure 4, we can see that as the neighborhood size $K$ increases, there is a gradual decline in precision. However, if $K$ is too small, the exploration is insufficient, which also reduces accuracy, as shown in the experiments with Llama-3.1-8B. Therefore, balancing $K$ is crucial to ensure both broad enough exploration and sufficient approximation accuracy.
Figure 4: Impact of different Neighborhood Size $K$ on GlitchMiner’s performance
Figure 5: Impact of different Batch Size $B$ on GlitchMiner’s performance
Effect of Batch Size. As shown in figure 5, different models exhibit varying sensitivity to batch size $B$. For example, Llama-3.1-8B and Qwen2.5-7B perform best at $B=4$, while Phi-3-mini peaks at $B=8$. Gemma-2-2B, on the other hand, reaches its highest performance at $B=16$ and maintains strong results at larger $B$ values. This indicates that the optimal $B$ value depends on the specific model architecture.
4.4 Token Entropy Analysis
To further validate the effectiveness of our entropy-based approach in detecting glitch tokens, we conducted an entropy analysis comparing glitch tokens and normal tokens across different models. For each model, we computed the average entropy of glitch tokens ($E_{\text{Glitch}}$) and normal tokens ($E_{\text{Normal}}$).
Figure 6 presents the comparison of average entropy values between glitch tokens and normal tokens for each evaluated model. As shown in the figure, glitch tokens consistently exhibit significantly higher entropy than normal tokens across all models.
Figure 6: Average entropy comparison between glitch tokens and normal tokens across different models. Glitch tokens have higher entropy, indicating greater uncertainty in the model’s predictions for these tokens.
This pronounced difference in entropy values indicates that models are more uncertain when predicting glitch tokens compared to normal tokens. The higher entropy of glitch tokens validates our hypothesis that maximizing entropy effectively guides the search towards tokens that are challenging for the model to predict.
Moreover, the consistent pattern of higher entropy for glitch tokens across diverse model families—including Llama, Qwen, Gemma, Phi-3, and Mistral—demonstrates the generality and robustness of our entropy-based approach. This suggests that our method can be effectively applied to a wide range of LLMs with different architectures and tokenization strategies.
These findings reinforce the effectiveness of GlitchMiner’s entropy-based optimization in efficiently detecting glitch tokens by focusing on areas of high prediction uncertainty within the model.
5 Conclusion
In conclusion, we introduce GlitchMiner, a novel method for efficiently mining glitch tokens in large language models through gradient-based discrete optimization. By leveraging entropy as the loss function and integrating a local search strategy, GlitchMiner accurately identifies tokens that deviate from normal distributions without relying on manually defined priors. Extensive experimental results demonstrate that GlitchMiner outperforms existing state-of-the-art methods across various model architectures, showcasing higher precision and adaptability. Ablation studies further confirm the contributions of key components to the overall performance. The consistent results across different models emphasize the robustness and generalizability of GlitchMiner. Future work may explore applying this method to other high-dimensional tasks that require efficient discrete optimization.