### v4

# GlitchMiner: Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety. Existing detection methods rely on predefined patterns, limiting their adaptability across diverse LLM architectures. We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by introducing entropy as a measure of prediction uncertainty and employing a local search strategy to explore the token space. Experiments across multiple LLM architectures demonstrate that GlitchMiner outperforms existing methods in detection accuracy and adaptability, achieving over 10% average efficiency improvement. This method enhances vulnerability assessment in LLMs, contributing to the development of more robust and reliable applications. Code is available at https://github.com/wooozihui/GlitchMiner.
1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 . However, with the growing reliance on these models in critical domains, a subtle yet significant vulnerability has emerged: glitch tokens geiping2024coercing . These anomalous tokens can severely disrupt an LLM’s output, leading to repetitive errors, nonsensical responses, or even harmful content. Such disruptions are particularly concerning in high-stakes applications where model reliability is crucial.
Recent research has made progress in detecting glitch tokens. Methods such as Magikarp Magikarp and GlitchHunter GlitchHunter focus on abnormal embedding patterns and clustering properties. However, these approaches are limited by their dependence on predefined patterns, restricting their adaptability across different LLM architectures and potentially missing novel forms of glitch tokens that deviate from known patterns.
In this paper, we introduce GlitchMiner, a novel framework for glitch token detection that overcomes the limitations of pattern-based heuristics by employing gradient-based discrete optimization to directly analyze a model’s prediction behavior. Central to our method is the use of entropy as an exploratory metric, guiding the search through the token space to identify regions indicative of unexpected model behavior.
The key innovations of GlitchMiner include:
  • Entropy-Based Optimization: By introducing entropy as a novel objective for glitch token detection, we systematically identify tokens that induce high uncertainty in model predictions. This approach enables detection of a broader range of glitch tokens, extending beyond traditional pattern- or embedding-based methods.
  • Gradient-Guided Local Search: Combining first-order Taylor approximation with a local search strategy, GlitchMiner efficiently explores the vast token space. This significantly enhances both the efficiency and accuracy of glitch token detection, allowing thorough exploration while maintaining computational feasibility.
  • Adaptive Cross-Architecture Detection: GlitchMiner is designed to adapt to the unique characteristics of various LLM architectures without relying on model-specific assumptions, enabling effective glitch token detection across a wide range of models.
We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families. The results show that GlitchMiner surpasses existing methods in detection efficiency in 90% of cases. Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1, with respective improvements of 10.5% and 10.7% over state-of-the-art methods. These gains enhance the detection of potentially harmful tokens, improving both accuracy and query efficiency. Such advancements are essential for LLM reliability, especially in applications where undetected glitch tokens could have serious consequences.
The remainder of this paper is structured as follows: Section 2 provides a review of related work in glitch token detection and gradient-based optimization techniques. Section 3 details the design principles and implementation of GlitchMiner. Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures. Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research.
2 Background and Related Work
2.1 What is a Glitch Token?
Glitch tokens, a concept that gained attention following the discovery of the "SolidGoldMagikarp" phenomenon by Rumbelow et al. solidgoldmagikarp , are anomalous tokens in LLMs that can trigger unexpected and often erroneous behaviors when processed by the model. These tokens typically result from irregularities in the training process, such as underrepresentation in the training data or inconsistencies in tokenization Magikarp ; geiping2024coercing . Glitch tokens have been found to exhibit several key characteristics:
  • They often have abnormal embedding patterns, such as unusually small $\ell_{2}$ norms or atypical positions in the embedding space Magikarp ; solidgoldmagikarp .
  • When input to an LLM, they can cause the model to produce repetitive, nonsensical, or completely unrelated outputs solidgoldmagikarp ; GlitchHunter .
  • They may lead to significant deviations in the model’s internal activations compared to normal tokens GlitchProber .
  • Glitch tokens can sometimes bypass content filters or trigger unexpected model behaviors, potentially compromising the safety and reliability of LLMs geiping2024coercing .
The study of glitch tokens is crucial for understanding and mitigating potential vulnerabilities in LLMs, as these tokens can potentially be exploited for adversarial attacks or lead to unintended model behaviors in critical applications geiping2024coercing .
2.2 Glitch Token Detection
The detection of glitch tokens in LLMs has become an increasingly important area of research, given their potential to disrupt model performance and reliability. The fundamental approach to identifying these anomalous tokens often relies on repetition tasks, a method popularized by the SolidGoldMagikarp study solidgoldmagikarp . In these tasks, the model is prompted to repeat a given token, with the assumption that normal tokens should be easily reproducible, while glitch tokens often lead to failures in this simple task.
Building upon this basic principle, researchers have developed several sophisticated methods to detect glitch tokens more efficiently and accurately:
Magikarp solidgoldmagikarp pioneered a fast and lightweight approach by analyzing token embeddings. It identifies potentially problematic tokens by examining characteristic patterns in the embedding space, such as unusual $\ell_{2}$ norms. These candidate tokens are then verified using repetition tasks, allowing for quick detection through direct queries to the model.
GlitchHunter GlitchHunter introduced a clustering-based method. This approach is founded on the observation that glitch tokens often cluster together in the embedding space. GlitchHunter constructs a Token Embedding Graph (TEG) to represent token relationships and applies the Leiden algorithm traag2019louvain to identify potential glitch clusters. These clusters are then refined through iterative hypothesis testing to improve detection accuracy.
GlitchProber GlitchProber takes a different approach by analyzing the internal activations within transformer layers, such as attention heads and hidden states. It reduces the dimensionality of these activations using PCA PCA and applies SVM classifiers SVM to identify glitch tokens. Additionally, it integrates mitigation mechanisms by modifying neuron activations to minimize the impact of glitch tokens during inference.
While these methods have significantly advanced our ability to detect glitch tokens, they are inherently limited by their reliance on pattern-based heuristics. Magikarp and GlitchHunter focus on specific embedding patterns or clustering behaviors, while GlitchProber examines predefined patterns in activation spaces. This dependence on pattern-based approaches can constrain their adaptability across diverse LLM architectures and potentially overlook novel forms of glitch tokens.
Our proposed method, GlitchMiner, aims to address these limitations by employing gradient-based discrete optimization to directly analyze an LLM’s prediction behavior, potentially offering a more flexible and comprehensive approach to glitch token detection.
2.3 Gradient-based Discrete Optimization
Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG ; PEZ leverage gradient information to predict how individual tokens impact the loss function. These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
HotFlip hotflip uses the one-hot encoding of tokens to compute gradients and selects the token with the largest negative gradient to replace the current token, aiming to minimize the loss. However, it only evaluates one candidate token per iteration, which can lead to suboptimal predictions and reduced accuracy.
AutoPrompt autoprompt improves upon HotFlip by evaluating multiple candidate tokens in each iteration. Instead of relying on gradients from one-hot encodings, it utilizes token embedding gradients for loss estimation, enhancing prediction accuracy by considering a broader range of potential token replacements.
GCG GCG extends HotFlip by incorporating multi-candidate token selection, similar to AutoPrompt, but it still uses the one-hot encoding of tokens to compute gradients for loss estimation. Notably, GCG has been applied to automated jailbreaks shen2023anything in LLMs, efficiently searching for adversarial suffixes.
AutoPrompt and GCG both rely on large batch sampling to mitigate inaccuracies in gradient prediction. We identified that these inaccuracies arise from the inaccuracy of Taylor expansions when input tokens are distant from the original points. This overlooks a fundamental condition of Taylor approximation: its accuracy is highest for points close to the reference point.
Building on these works, we introduce a local search strategy in our approach. This improvement enables us to achieve high precision in gradient estimation without relying on large batch sampling, by focusing on a smaller, localized token space. By addressing the core issue of Taylor approximation accuracy, our method allows for more efficient and accurate exploration of the token space, which is particularly valuable for glitch token detection.
3 Method
Algorithm 1 GlitchMiner Framework
1:Input: Token set $\mathcal{T}$, Iteration number $I$, Batch size $B$
2:Output: Glitch token set $\mathcal{G}$
3:# Stage 1: Initialization
4:$\mathcal{T}^{*}\leftarrow\text{Filter}(\mathcal{T})$ Filter out the unnecessary token set $\mathcal{T}^{*}$
5:$\mathcal{G}\leftarrow\emptyset$
6:# Stage 2: Mining
7:for $i\in I$ do
8:     Select a batch of tokens $\mathcal{B}$ from $\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where $|\mathcal{B}|=B$
9:     for each token $t\in\mathcal{B}$ do
10:         if Verify($t$then
11:              Add $t$ to $\mathcal{G}$
12:         else
13:              Add $t$ to $\mathcal{T}^{*}$
14:         end if
15:     end for
16:end for
GlitchMiner Framework. As shown in Algorithm 1, GlitchMiner operates in two main stages: initialization and mining.
In the initialization stage, we filter out tokens that don’t require detection to reduce the search space and improve efficiency, after which the mining stage iteratively selects batches of unverified tokens for verification. Each token is checked: if it is identified as a glitch token, it is added to the discovered set; otherwise, it is marked unnecessary for further detection. This process continues until a stopping condition is met, such as reaching a specified number of iterations or detected glitch tokens.
The following sections detail how GlitchMiner implements its key components:
  1. 1.
    Token Filter: Identifying tokens that don’t require detection.
  2. 2.
    Glitch Token Verification: Methods to verify if a token is a glitch token.
  3. 3.
    Token Selection Strategy: Selecting the most promising glitch token candidates in each iteration.
By exploring these components, we demonstrate how GlitchMiner efficiently implements glitch token mining in practice.
3.1 Token Filter
In the initialization stage of GlitchMiner, we first filter all tokens to remove those that do not need to be detected. Similar to Magikarp , we designed a token filter module that classifies tokens based on their characteristics and filters out those that are unnecessary for detection.
The core idea of the token filter module is to classify each token by decoding and then re-encoding it, ensuring it meets specific classification criteria. Specifically, we prepend a special prefix "«" to each token to maintain consistency during the encoding and decoding process. Then, we filter out tokens that fall into the following categories:
  • SPECIAL: Special tokens, such as [BOS], </s>, etc., which are enclosed in brackets or angle brackets and contain alphabetic characters.
  • UNDECODEABLE: Tokens that cannot be decoded, usually containing illegal characters.
  • UNREACHABLE: Tokens that cannot be restored to their original token ID through the decoding and re-encoding process.
During the classification process, we first decode each token ID to obtain its corresponding string representation. If decoding fails, the token is classified as UNDECODEABLE. Next, we encode the decoded string and check if it can be restored to the original token ID. If it cannot, the token is classified as UNREACHABLE. If it meets the characteristics of a special token, it is classified as SPECIAL.
By filtering out these unnecessary tokens, we retain only the essential ones for subsequent mining, which improves the algorithm’s efficiency and ensures the accuracy of the detection results.
3.2 Glitch Token Verification
Our glitch token verification process consists of two stages: optimization-driven detection and cross-verification.
Template for Optimization. For our primary detection and optimization process, we designed the following task template:
User: Please repeat the string: "«{token}»"
Assistant: Sure, the string is: "«{token}
In this template, the blue text represents the prompt, the green text {token} is the input token being tested, and the red text {token} represents the model’s predicted output. If the predicted token does not match the input token, it is initially classified as a potential glitch token.
This template is specifically designed to support our gradient-based optimization process. Its simplicity allows for clear gradient computations, while the prefilled assistant response ensures a uniform starting point for each token evaluation. Wrapping the token in "«" and "»" symbols minimizes interference from surrounding context. These features enable precise entropy calculations and gradient estimations, which are crucial for our entropy-guided search strategy described in the token selection section.
Cross-Verification. While our primary template is optimized for the search process, relying solely on one template may lead to false positives. To enhance the robustness of our detection, we implement a cross-verification step using two additional templates derived from GlitchHunter and Magikarp. Each potential glitch token identified by our primary template undergoes verification with these additional templates. A token is confirmed as a glitch token only if it fails the repetition task across all templates.
This two-stage approach combines the efficiency of our optimization-driven search with the reliability of multi-template verification, significantly reducing false positives while maintaining the effectiveness of our gradient-based detection method.
3.3 Token Selection
Motivation. Previous methods for glitch token detection often rely on manually observed embedding patterns. However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens. To address these limitations, we propose an entropy-based approach. Entropy, as a measure of uncertainty in probability distributions, offers a model-agnostic way to identify tokens that cause unexpected behavior in LLMs. By focusing on tokens that maximize entropy, we can detect glitch tokens that deviate significantly from normal token behavior, regardless of their specific characteristics or the underlying model architecture.
Entropy-Guided Exploration. We define the entropy $H(t)$ for a token $t$ as:
$H(t)=-\sum_{v\in\mathcal{V}}P(v\mid\mathbf{h}(t))\log P(v\mid\mathbf{h}(t))$
where $\mathbf{h}(t)$ is the context embedding for token $t$, and $P(v\mid\mathbf{h}(t))$ is the model’s predicted probability distribution over the vocabulary $\mathcal{V}$. By maximizing entropy, we aim to find tokens that cause high uncertainty in the model’s predictions, potentially indicating glitch behavior.
Optimization Objective. Our goal is to find a batch of tokens $\mathcal{B}$ that maximizes the total entropy:
$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{T}_{c},|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}H(t)$
Here, $\mathcal{T}_{c}$ represents the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where:
  • $\mathcal{T}$ is the initial set of all tokens in the model’s vocabulary.
  • $\mathcal{T}^{*}$ is the set of tokens that have been filtered out or verified as non-glitch tokens. This includes special tokens, undecoded tokens, and tokens that have been checked and confirmed as normal.
  • $\mathcal{G}$ is the set of identified glitch tokens discovered so far in the process.
By excluding $\mathcal{T}^{*}$ and $\mathcal{G}$ from $\mathcal{T}$, we ensure that our search focuses only on the remaining unverified tokens, improving efficiency and preventing redundant checks.
Local Search Strategy. To efficiently solve this optimization problem while addressing the limitations of global Taylor approximation, we introduce a local search strategy. We begin by selecting an initial token $t_{0}$ and computing its entropy $H(t_{0})$ and gradient $\nabla_{e}H(t_{0})$. We then define a local neighborhood $\mathcal{N}_{K}(t_{0})$ consisting of the $K$ nearest neighbors to $t_{0}$ in the embedding space. This localized approach allows us to refine the search within a region close to $t_{0}$, improving approximation accuracy and ensuring efficient exploration of potential glitch tokens.
Within this local neighborhood, we estimate the entropy of each candidate token $t\in\mathcal{N}_{K}(t_{0})$ using first-order Taylor approximation:
$\hat{H}(t)\approx H(t_{0})+\nabla_{e}H(t_{0})^{T}(e_{t}-e_{t_{0}})$
where $e_{t}$ and $e_{t_{0}}$ are the embedding vectors of tokens $t$ and $t_{0}$, respectively. This approximation allows us to efficiently estimate the entropy change for nearby tokens without the computational cost of exact calculations.
Based on these estimates, we select a batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy. For this selected batch, we then compute the actual entropy values and choose the token with the highest entropy as the starting point for the next iteration. This iterative process ensures that our search is guided by the most promising candidates at each step.
This local search strategy significantly improves the accuracy of entropy estimation by focusing on tokens close to the reference point, addressing a key limitation of previous gradient-based methods. By balancing exploration of the token space with exploitation of local information, our approach enables more efficient and effective glitch token detection.
The combination of entropy-guided exploration and local search allows our method to adapt to diverse LLM architectures without relying on model-specific assumptions. This results in a more robust and versatile approach to glitch token detection, capable of identifying a wide range of anomalous tokens across different model types.
4 Experiments
4.1 Experimental Setup
Test LLMs. We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach. The selected models include Meta’s Llama series llama2 ; llama3_1 , Alibaba’s Qwen models qwen2 ; qwen2_5 , Google’s Gemma models gemma2 , Microsoft’s Phi-3 models phi3 , and Mistral models mistral_7B ; mistral2024nemo . The details are presented in Table 1.
Model Family Model Names
Llama Models Llama-3.1-8B-Instruct , Llama-2-7B-chat-hf
Qwen Models Qwen2.5-7B-Instruct , Qwen2-7B-Instruct
Gemma Models Gemma-2-2b-it, Gemma-2-9b-it
Phi-3 Models Phi-3-mini-128k-instruct, Phi-3.5-mini-instruct
Mistral Models Mistral-7B-Instruct-v0.3, Mistral-Nemo-Instruct-2407
Table 1: Test LLMs used in the experiments.
Evaluation Metrics. To assess the performance of our glitch token detection method, we employ the Detected@N metric. This metric measures the absolute number of true glitch tokens identified within the top $N$ predictions made by the algorithm. It is calculated as:
[…]
For example, Detected@1000 measures the number of true glitch tokens discovered within the top 1000 predicted tokens. This metric directly reflects a method’s effectiveness in identifying glitch tokens under practical constraints, as it captures both the detection capability and query efficiency of the approach. By comparing the number of glitch tokens detected under the same query budget ($N$), we can effectively evaluate different methods’ ability to maximize token discovery while minimizing computational resources. This makes it an ideal metric for real-world applications where efficient detection of glitch tokens is crucial.
Baselines. We compare our proposed glitch token detection method with two state-of-the-art approaches: GlitchHunter GlitchHunter and Magikarp Magikarp . These methods serve as the primary benchmarks for evaluating our approach.
Although GlitchProber GlitchProber is another relevant method, it follows a fundamentally different approach by pre-collecting a subset of glitch tokens to train a classifier, introducing a supervised learning component. In contrast, GlitchMiner, along with GlitchHunter and Magikarp, uses heuristic-based methods to detect glitch tokens without relying on labeled data or additional classifier training. This methodological difference makes a direct comparison less meaningful, so we focus our evaluation on methods that align more closely with our unsupervised approach.
Parameter Settings. In our implementation of GlitchMiner, we use $K$=32 and $B$=8 as the default parameters. These values were chosen based on empirical testing to balance computational efficiency and detection effectiveness. Specifically, $K$=32 defines the size of the local neighborhood considered in each iteration, while $B$=8 determines the batch size for entropy computation. These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
Initialization Strategy in Experiments. To ensure stable and consistent comparisons across runs, we initialize the search with the token exhibiting the smallest $\ell_{2}$ norm in the embedding space, based on prior observations that such tokens often exhibit glitch-like behaviors. However, as shown in Figure 4, we found that GlitchMiner remains robust to different initialization choices, achieving similar performance even with random starting points.
4.2 Main Results
Model Metric GlitchHunter Magikarp GlitchMiner (ours)
Llama-3.1-8B-Instruct Detected@1000 25 664 568
Detected@2000 56 935 1164
Llama-2-7B-chat-hf Detected@1000 61 100 319
Detected@2000 126 186 532
Qwen2.5-7B-Instruct Detected@1000 75 1000 1000
Detected@2000 180 1893 1839
Qwen2-7B-Instruct Detected@1000 96 999 1000
Detected@2000 191 1842 1847
Gemma-2-2b-it Detected@1000 23 678 744
Detected@2000 35 984 1019
Gemma-2-9b-it Detected@1000 29 623 775
Detected@2000 45 983 1089
Phi-3.5-mini-instruct Detected@1000 20 393 396
Detected@2000 44 496 516
Phi-3-mini-128k-instruct Detected@1000 26 398 404
Detected@2000 55 489 517
Mistral-7B-Instruct-v0.3 Detected@1000 6 110 219
Detected@2000 19 130 302
Mistral-Nemo-Instruct-2407 Detected@1000 48 574 695
Detected@2000 79 918 976
Average Detected@1000 40.9 553.9 612.0
Detected@2000 93.0 885.6 980.1
Table 2: Detected@1000 and Detected@2000 comparison of methods across different models.
Table 2 displays the performance comparison of GlitchMiner with the state-of-the-art methods GlitchHunter and Magikarp across different LLM architectures using the Detected@N metric. The results highlight GlitchMiner’s strong and versatile detection capabilities.
In the majority of cases, GlitchMiner achieved the highest Detected@2000 score, consistently outperforming both baselines in terms of identified glitch tokens within the top 2000 predictions. This is particularly evident in models such as Llama-2-7B-chat-hf, where GlitchMiner achieved a Detected@2000 of 532, surpassing Magikarp by a notable margin. For the Qwen and other models, GlitchMiner maintained robust performance, demonstrating adaptability across different LLM architectures.
Figure 1: Comparison of GlitchMiner performance with and without local search strategy
These experimental findings underscore GlitchMiner’s ability not only to accurately detect glitch tokens but also to generalize across a diverse set of model architectures. This adaptability and precision position GlitchMiner as a powerful tool for enhancing the robustness and security of LLMs.
4.3 Ablation Study
To evaluate the contributions of key components in GlitchMiner, we conducted ablation studies focusing on the local search strategy, neighborhood size $K$, batch size $B$, and initialization token.
Effect of Local Search. The local search strategy plays a crucial role in enhancing GlitchMiner’s ability to detect glitch tokens by improving the precision of the Taylor approximation. Without local search, detection accuracy drops significantly (Figure 1), as global search lacks the necessary granularity to maintain precise approximations within the token space.
Effect of Neighborhood Size. We analyzed the impact of neighborhood size $K$ on detection performance. As shown in Figure 2, increasing $K$ generally leads to a decline in Detected@1000 values across models. This trend indicates that as $K$ grows, the Taylor approximation becomes less effective, resulting in reduced prediction accuracy.
Figure 2: Impact of different Neighborhood Size $K$ on GlitchMiner’s performance
Figure 3: Impact of different Batch Size $B$ on GlitchMiner’s performance
Figure 4: Effect of Initialization Method on GlitchMiner’s Detected@1000 score.
Effect of Batch Size. As shown in Figure 3, the performance of GlitchMiner remains relatively stable as batch size $B$ increases. Notably, even with $B=1$, GlitchMiner achieves effective detection results, indicating that it can make accurate predictions without relying on a large batch size.
Effect of Initialization Token. As shown in Figure 4, GlitchMiner’s performance remains stable across different initialization tokens. The red dots represent the minimum $\ell_{2}$ norm initialization, while the orange dots show three random trials. For most models, random initialization results are close to the minimum $\ell_{2}$ norm, indicating that GlitchMiner achieves consistent detection accuracy regardless of the initialization approach.
4.4 Token Entropy Analysis
To further validate the effectiveness of our entropy-based approach in detecting glitch tokens, we conducted an entropy analysis comparing glitch tokens and normal tokens across different models. For each model, we computed the average entropy of glitch tokens ($E_{\text{Glitch}}$) and normal tokens ($E_{\text{Normal}}$).
Figure 5 presents the comparison of average entropy values between glitch tokens and normal tokens for each evaluated model. As shown in the figure, glitch tokens consistently exhibit significantly higher entropy than normal tokens across all models.
Figure 5: Average entropy comparison between glitch tokens and normal tokens across different models. Glitch tokens have higher entropy, indicating greater uncertainty in the model’s predictions for these tokens.
This pronounced difference in entropy values indicates that models are more uncertain when predicting glitch tokens compared to normal tokens. The higher entropy of glitch tokens validates our hypothesis that maximizing entropy effectively guides the search towards tokens that are challenging for the model to predict.
Moreover, the consistent pattern of higher entropy for glitch tokens across diverse model families—including Llama, Qwen, Gemma, Phi-3, and Mistral—demonstrates the generality and robustness of our entropy-based approach. This suggests that our method can be effectively applied to a wide range of LLMs with different architectures and tokenization strategies.
These findings reinforce the effectiveness of GlitchMiner’s entropy-based optimization in efficiently detecting glitch tokens by focusing on areas of high prediction uncertainty within the model.
5 Conclusion
In this paper, we introduced GlitchMiner, a novel framework for detecting glitch tokens in LLMs through gradient-based discrete optimization. Our method combines entropy-based loss functions with a local search strategy to navigate token space efficiently, improving both detection accuracy and computational efficiency. Experimental results across diverse LLMs highlight GlitchMiner’s robustness and versatility compared to existing methods. Our findings confirm GlitchMiner’s effectiveness in identifying glitch tokens across architectures, demonstrating its potential for enhancing the security and reliability of LLM-based applications. Future work could explore extending this approach to other high-dimensional discrete optimization tasks.
Broader Impact
The potential impact of this research extends beyond mere academic interest. By improving the detection of glitch tokens, GlitchMiner contributes to enhancing the overall reliability and safety of LLMs in real-world applications. This is particularly crucial in high-stakes domains such as healthcare, finance, and autonomous systems, where the consequences of model errors can be severe. Furthermore, our approach opens new avenues for understanding and mitigating vulnerabilities in AI systems, potentially influencing the development of more robust and trustworthy AI technologies in the future.
Acknowledgments
We would like to express our sincere gratitude to Sander Land for his valuable insights and feedback on this work, particularly regarding the evaluation metrics. His suggestions were instrumental in refining our approach and improving the clarity of our results. We also wish to thank Guanlin Li for his valuable suggestions on our manuscript. His insights were crucial for improving the content and structure of the paper.